# VersusDebias: Universal Zero-Shot Debiasing for Text-to-Image Models via SLM-Based Prompt Engineering and Generative Adversary

### v2

@@ Abstract
With the rapid development of Text-to-Image models, biases in human image generation against demographic groups social attract more and more concerns. Existing methods are designed based on certain models with fixed prompts, unable to accommodate the trend of high-speed updating of Text-to-Image (T2I) models and variable prompts in practical scenes. Additionally, they fail to consider the possibility of hallucinations, leading to deviations between expected and actual results. To address this issue, we introduce VersusDebias, a novel and universal debiasing framework for biases in T2I models, consisting of one generative adversarial mechanism (GAM) and one debiasing generation mechanism using a small language model (SLM). The self-adaptive GAM generates specialized attribute arrays for each prompts for diminishing the influence of hallucinations from T2I models. The SLM uses prompt engineering to generate debiased prompts for the T2I model, providing zero-shot debiasing ability and custom optimization for different models. Extensive experiments demonstrate VersusDebias’s capability to rectify biases on arbitrary models across multiple protected attributes simultaneously, including gender, race, and age. Furthermore, VersusDebias outperforms existing methods in both zero-shot and few-shot situations, illustrating its extraordinary utility. Our work is openly accessible to the research community to ensure the reproducibility.
@@ 1 Introduction
Text-to-Image (T2I) models, as a crucial part of Artificial Intelligence Generated Content (AIGC) technology, are evolving at an extraordinary speed due to the low barriers to entry and open-source initiatives by companies such as Stability AI (Rombach et al. 2022; Luo et al. 2023; Betker et al. 2023). Many models are able to generate high-quality and photo-realistic images from simple textual prompts. Nevertheless, similar to other generative models (Ross, Katz, and Barbu 2020; Mehrabi et al. 2021; Zhou et al. 2024), the growing adoption also raised concerns about fairness and biases in generated images, especially in images depicting human (Bansal et al. 2022; Cho, Zala, and Bansal 2023; Luccioni et al. 2023). Some research shows that the models have obvious biases to specific social groups (Perera and Patel 2023), especially women and colored races.
Some researchers have proposed their solutions to address the issue (Gandikota et al. 2024; Schramowski et al. 2023; Shrestha et al. 2024). However, these methods face similar problems. Neither of them solves the two most important problems that stop debiasing methods to practical application. First, all existing methods are only applicable to a certain T2I model, usually SD 1.5 (Rombach et al. 2022). However, with the widely application of distillation (Meng et al. 2023) and the participation of communities like HuggingFace, T2I models evolve rapidly. For example, the mainstream base model in the community evolves from SD1.5 to SDXL to SDXL Lightning (Lin, Wang, and Yang 2024) in only one year, and cutting-edge developers proposes models with new architectures such as Stable Cascade (Pernias et al. 2023) and Pixart (Chen et al. 2023), but recent research about debiasing still develops their models based on early models, limiting the application prospects. Second, most methods considers only specific debiasing metrics and are unable to zero-shot debias. For example, Fair Diffusion (Friedrich et al. 2023) uses a look-up table to add protected attributes to certain prompts about occupations in the table to ensuring demographic fairness. However, this method fails to tackle with any prompts that are not in the table. Debiasing techniques require good robustness and the ability of zero-shot generation to be adaptive for various inputs.
Figure 1: VersusDebias decreases biases in generated images by prompt engineering via the executor (a finetuned SLM).
@@
To address this, we propose VersusDebias, a universal zero-shot debiasing framework for T2I models, consisting of a generative adversarial mechanism (GAM) and a debiasing generation mechanism (DGM). A simple workflow is shown in Figure 1. Similar to generative adversarial networks (GAN) (Goodfellow et al. 2020), our GAM consists of a generator and a discriminator. This component does not directly participate in the final image generation but provides a dataset for the executor to modify the prompts. Through adversarial generation, GAM produces debiased attribute arrays corresponding to 2,060 prompts, which consist of protected attributes from the ignored social groups. The main body of our DGM is an fine-tuned small language model (SLM), which ensures performance while significantly reducing computational resource consumption compared to traditional LLMs. The SLM can accurately identify potentially biased attributes in complex prompts and contextually add protected attributes according to debiased attribute arrays to generate debiased prompts, providing zero-shot debiasing capability. Since the generator in GAM and the T2I model used in DGM are the same, debiased attribute arrays can mitigate the hallucinations of the T2I model through post-processing. All models in VersusDebias can adapt to a new model through inference alone without retraining, which circumvents the computational overhead. Additionally, VersusDebias is designed as a universal plug-in module, equipped with interfaces based on ComfyUI API (Comfyanonymous 2024) and Websocket. This allows it to adapt to thousands of models compatible with ComfyUI simply by replacing the workflow file. This structure, combined with the adaptive capabilities provided by GAM, enables VersusDebias to reduce biases in any T2I model.
@@
To evaluate the effectiveness of our work, we compare VersusDebias with other methods from two perspectives. First, we evaluate the demographic bias of three standard models before and after incorporating VersusDebias. Second, we compare VersusDebias with existing debiasing methods. All comparisons are conducted using both small-scale human evaluations and large-scale machine evaluations based on multi-modal large language model (MLLM), considering both zero-shot and few-shot scenarios. The results demonstrate that VersusDebias effectively reduces biases in T2I models and significantly outperforms other methods.
@@
  • We propose VersusDebias, a novel and universal framework to improve demographic fairness and decrease biases in T2I models which consists of a GAM and a DGM.
  • VersusDebias integrates numerous advanced technologies, including MLLM, SLM, prompt engineering, and generative adversarial algorithms.
  • VersusDebias implements zero-shot debiasing with easily transferable ComfyUI API interfaces, which not only demonstrates its universality but also significantly lowers the barriers to using debiasing technology, advancing its practical application.
@@ 2 Related Works
Over the last few years, T2I models like Stable Diffusion and DALL-E have seen increasing adoption (Ramesh et al. 2022; Esser et al. 2024). Many T2I models can generate high-quality images efficiently. Nonetheless, several studies prove that these models inherit biases from their training da tasets (Chinchure et al. 2023; Wan et al. 2024). Recent research classifies these biases into implicit generative bias and explicit generative bias (Luo et al. 2024a), corresponding to the sociological concepts of implicit bias (Pritlove et al. 2019) and explicit bias (Fridell 2013). Implicit generative bias refers to the phenomenon where, without specific instructions on protected attributes, T2I models tend to generate images that do not consist of demographic realities. For instance, when asked to generate ”a photo of a surgeon”, models tend to produce images featuring male surgeons. Explicit generative bias is a specific type of hallucination, referring to the phenomenon where T2I models tend to generate images that do not consist with prompts containing specific protected attributes. For example, when asked to generate ”a photo of a rich black person” models may sometimes fail to correctly generate an image featuring a black person. In our work, we primarily address the issue of implicit generative bias and avoid the influence of explicit generative bias which disturbs conventional methods.
@@
Recent works attempts to debias T2I models at different levels, such as diffusion-level (Shrestha et al. 2024) and prompt-level (Clemmer, Ding, and Feng 2024). Diffusion-level methods interfere with the image generation process, which can lead to uncontrollable side effects on the quality and style of the final image. Moreover, these methods can only be applied to diffusion-based models and are not suitable for transformer-based models like DALL-E3 (Betker et al. 2023). In contrast, debiasing based on prompt engineering, as a post-processing method, is more adjustable and more broadly applicable with less impact on image quality. However, these methods also face serious challenges. First, traditional prompt engineering methods rely on the assumption that the model correctly generate images based on the prompt, which cannot avoid hallucinations, especially those caused by explicit generative bias. Second, prompt engineering might have unintended side effects on unrelated attributes. For example, using SDXL-Turbo (Sauer et al. 2023), the proportion of women generated with the prompt ”a photo of a South Asian tennis player” is significantly lower than that generated with the prompt ”a photo of a tennis player” (Luo et al. 2024b). Lastly, the severity of these issues varies across different models, meaning that traditional methods, even if they address the first two issues, can only be applied to specific models, greatly limiting their practicality. In contrast, our method employs an innovative GAM structure that effectively addresses these problems, providing unprecedented practicality.
Figure 2: Pipeline of VersusDebias at top-level. The yellow boxes represent the components of the Discriminator. The three gray rectangles distinguish the three main stages of debiasing. The blue rectangles represent data that change with the T2I model, while the red rectangles represent data that is independent of the T2I model.
@@
Due to the tremendous success of ChatGPT-4 (Achiam et al. 2023), LLMs, especially MLLMs, have garnered significant attention and are rapidly evolving. The improved accuracy of MLLMs addresses the significant shortcomings of traditional alignment models like CLIP (Radford et al. 2021) and BLIP-2 (Li et al. 2023) in recognizing race and age (Luo et al. 2024b; Chinchure et al. 2023), providing essential support for the implementation of the discriminator in GAM. On the other hand, the development of edge AI attracts more researchers to focus on SLMs (Bai et al. 2023; Zhang et al. 2024). These SLMs retain most of the original performance of LLMs while significantly reducing computational resource consumption. For prompt engineering, compared to traditional methods, which involve retrieving and replacing keywords from a predefined set, LLM-based methods exhibit higher precision and zero-shot generation capabilities (Clemmer, Ding, and Feng 2024). In VersusDebias, we utilize a recent SLM and a MLLM to achieve outstanding debiasing performance and accuracy with minimal computational resources.
@@ 3.1 Overview
In this section, we introduce the top-level pipeline of VersusDebias. A brief overview of the pipeline is shown in Figure 2. As illustrated, VersusDebias consists of a GAM, a DGM, and the corresponding dataset.
To debias a model, VersusDebias first uses prompts without protected attributes from the dataset to drive the generator to generate images, which are then conveyed into the discriminator. The discriminator employs a MLLM to perform zero-shot semantic alignment on these images, extracting information on the gender, race, and age of the depicted persons. The alignment results are statistically averaged to determine the generative demographic proportions of three protected attributes for each main prompt, the occupation. The discriminator compares the proportions of each gender, race, and age group to the ground truth and modifies the prompts by adding the underrepresented protected attributes, which ensures that the generator produces more images featuring these ignored social groups in the next iteration. Each main prompt corresponds to a protected attribute array. The newly generated images are then conveyed into the discriminator again, and the process is repeated until the specified number of epochs.
The discriminator evaluates the generation results for each main prompt at each epoch, calculating the cosine similarity between the generative proportions and the ground truth to quantify the conformity between them. The protected attribute array with the highest cosine similarity for each main prompt across different epochs is then combined into a single dictionary, referred to as the best results, and conveyed to the DGM as shown in Figure 2.
Next, the executor uses the fine-tuned SLM to accurately extract the potentially bias-inducing main prompt from the user prompt. It then selects an appropriate protected attribute based on the best results and adds it to the original prompt. Finally, the generator of the DGM uses the processed prompt to generate debiased images.
@@ 3.2 Generator
As a universal debiasing framework based on prompt engineering, VersusDebias can use any model in its generator, including diffusion-based models (Nichol et al. 2021; Sauer et al. 2023; Esser et al. 2024; Lin, Wang, and Yang 2024; Song, Sun, and Yin 2024) and transformer-based models (Dayma et al. 2021; Ding et al. 2022; Ramesh et al. 2022; Chen et al. 2023). To lower the barrier of VersusDebias and attract more researchers to debiasing studies, we develop an interface compatible with ComfyUI, which is a widely-used GUI for thousands of T2I models based on Stable Diffusion (Rombach et al. 2022). ComfyUI only requires changes to the input workflow file to adapt to various models and is equipped with a comprehensive API for external program calls. Additionally, due to its openness and flexibility, some models that use other architectures, such as Playground (Li et al. 2024), PixArt (Chen et al. 2024b), and Stable Cascade (Pernias et al. 2023), can be compatible with ComfyUI by simply installing specific extension nodes. For the reasons above, we choose ComfyUI as the tool to connect VersusDebias with the T2I models. In terms of communication, VersusDebias uses Websocket to monitor the status of ComfyUI and convey instructions. This approach not only achieves a fully automated debiasing process but also allows both systems to operate in completely independent environments, ensuring optimal compatibility.
In DGM, the generator uses the prompt processed by the executor to generate images based on user input. In GAM, the generator uses prompts from the dataset. The dataset creation process for this section is as follows. Based on the 607 occupations classified by the Bureau of Labor Statistics of the U.S. (BLS 2024a), we consolidated occupations that are typically not distinguished in everyday contexts, resulting in 103 occupations. We also added descriptions of image quality and character states with ChatGPT-4o (Achiam et al. 2023), generating 20 prompts for each occupation and totaling 2,060 prompts.
@@ 3.3 Discriminator
As the core part of the GAM, the discriminator consists of three components: the aligner, the array editor, and the evaluator. In the following sections, we provide a detailed explanation of the implementation of these three components.
@@
The function of the aligner is to extract information about the individuals in the images and compute the average of the extraction results for all prompts of the same main prompt. We used the fine-tuned InternVL-4B-1.5 (Chen et al. 2024c) from BIGbench in the aligner, which is trained via the dataset from FairFace (Karkkainen and Joo 2021) and reportedly achieves an alignment accuracy of 97.93%. We do a test at a sample of 50 images, whose result is shown in the supplementary material, indicating that this model exhibits high accuracy, meeting the requirements of VersusDebias.
@@
The function of the array editor is comparing the generative proportions to the ground truth and adding the underrepresented protected attributes to the prompt. To achieve this, the array editor generates a new dictionary by subtracting the ground truth values from the actual generative proportions. The keys in this dictionary are main prompts, with negative values indicating that the generated data is below the actual data and positive values indicating the opposite. For all protected attributes with negative differences, the array editor adds these underrepresented attributes to the attribute array. The length of the attribute array $L$ can be adjusted based on the requirements for speed and accuracy, but it should generally be maintained at 4 times the number of epochs. This is because there are a total of nine protected attributes, and assuming an equal distribution of positive and negative differences, approximately 4.5 attributes will be added each time. Allowing for a buffer, the multiplier is set to 4 as a floor value, and $L$ is at least 5 as a ceiling value. The array editor performs this operation once per epoch until the specified limit is reached. A detailed algorithm is shown in Algorithm 1.
@@
The ground truth used the same occupational classification as the prompt set. For age, we divided it into three stages: young, middle-aged, and elderly. For gender, due to the cognitive limitations of the language model, we classify gender into male and female. For race, we employ the four-category classification from FairFace (Karkkainen and Joo 2021), distinguishing race into White, Black, East Asian, and South Asian. For racial proportions, we used the global racial population data from the United Nations (UNDESA 2022). For gender and age, the proportions for practitioners for a occupation are calculated by dividing the number of individuals of a specific gender or age group by the total number of individuals within all detailed occupation belonging to the occupation, based on the data from the BLS (BLS 2024b, c).
@@
where $S_{k}$ is the cosine similarity for the protected attribute $k$ of a main prompt, $p_{k}i$ and $q_{k}i$ are the generative proportion and real demographic proportion of the $i$ attribute, and $n_{k}$ is the total number of the attributes for this protected attribute.
After calculating the cosine similarity for these three attributes separately, the evaluator computes a weighted average to obtain the cumulative cosine similarity. We set the weights for gender and race at 0.35 each and the weight for age at 0.3. This weighting reflects the general consensus that gender and racial discrimination issues are more prevalent and severe. The weights can be adjusted based on different debiasing requirements:
@@
where $S_{c}$ is the cumulative cosine similarity for a main prompt, $w_{k}$ is the weight for the protected attribute $k$.
Finally, the evaluator selects the best-performing debiased attribute arrays based on the cumulative cosine similarity and combines these arrays into a single dataset for subsequent fine-tuning of the SLM.

### v3

@@ Abstract
With the rapid development of Text-to-Image (T2I) models, biases in human image generation against demographic social groups become a significant concern, impacting fairness and ethical standards in AI. Some researchers propose their methods to tackle with the issue. However, existing methods are designed for specific models with fixed prompts, limiting their adaptability to the fast-evolving models and diverse practical scenarios. Moreover, they neglect the impact of hallucinations, leading to discrepancies between expected and actual results. To address these issues, we introduce VersusDebias, a novel and universal debiasing framework for biases in arbitrary T2I models, consisting of an array generation (AG) module and an image generation (IG) module. The self-adaptive AG module generates specialized attribute arrays to post-process hallucinations and debias multiple attributes simultaneously. The IG module employs a small language model to modify prompts according to the arrays and drives the T2I model to generate debiased images, enabling zero-shot debiasing. Extensive experiments demonstrate VersusDebias’s capability to debias any models across gender, race, and age simultaneously. In both zero-shot and few-shot scenarios, VersusDebias outperforms existing methods, showcasing its exceptional utility. Our work is accessible at https://github.com/VersusDebias/VersusDebias to ensure reproducibility and facilitate further research.
@@ 1 Introduction
Text-to-Image (T2I) models, as a crucial part of Artificial Intelligence Generated Content (AIGC) technology, are evolving at an extraordinary speed driven by open-source initiatives by companies such as Stability AI (Rombach et al. 2022; Luo et al. 2023; Betker et al. 2023), enabling even non-experts to create high-quality and photorealistic images from simple textual prompts. Nevertheless, similar to other generative models (Mehrabi et al. 2021; Zhou et al. 2024), the widespread adoption of T2I models has raised significant concerns regarding fairness and biases, particularly in images depicting human (Bansal et al. 2022; Luccioni et al. 2023). Existing research highlights systematic biases within T2I models to specific social groups. For instance, T2I models primarily depict respectable occupations like lawyers and CEOs as white middle-aged men (Cho, Zala, and Bansal 2023), ignoring women and colored races.
Some researchers have proposed their solutions to address the issue (Gandikota et al. 2024; Schramowski et al. 2023; Shrestha et al. 2024). However, none of these methods fully resolve the two most crucial challenges that hinder the practical application of debiasing techniques. First, existing approaches are typically tailored to specific T2I models, commonly SD 1.5 (Rombach et al. 2022), which limits their adaptability, especially given the rapid evolution of T2I models driven by the adoption of distillation (Meng et al. 2023) and the participation of communities. For instance, the primary base model iterates three times from SD1.5 to SDXL-L (Lin, Wang, and Yang 2024) within a year, while researchers introducing models with entirely new architectures (Pernias et al. 2023; Chen et al. 2023). Recent research still develops their methods based on outdated models, limiting their application prospects. Second, most methods lack zero-shot debiasing capabilities. For example, FairDiffusion (Friedrich et al. 2023) uses a predefined look-up table to recognize prompts and add gender description to improve fairness, failing to deal with prompts that fall outside the predefined table. Like other AIGC techniques, effective debiasing techniques need robustness and zero-shot debiasing abilities to handle diverse inputs, ensuring broader applicability.
Figure 1: A simple generation workflow of VersusDebias.
@@
To address this, we propose VersusDebias, a universal and self-adaptive zero-shot debiasing framework for T2I models, consisting of an array generation (AG) module and an image generation (IG) module. A simple workflow is shown in Figure 1. Referring the structure of generative adversarial networks (Goodfellow et al. 2020), our AG module employs a generator and a discriminator to perform generative adversary. The discriminator employs a multi-modal large language model (MLLM) to achieve exceptional alignment performance. The AG module produces debiased attribute arrays corresponding to 2,060 prompts, which consist of protected attributes from the ignored social groups, providing reference for modifying the prompts. Compared to existing methods that add protected attributes according to fixed demographic proportions, debiased attribute arrays mitigate the hallucinations of T2I models, i.e., instances where models generate images that are inconsistent with the prompt, through post-processing. The IG module employs a fine-tuned small language model (SLM), i.e., LLM with fewer parameters, which can accurately complete named entity recognition (NER) in complex prompts and contextually add protected attributes according to debiased attribute arrays to generate debiased prompts, providing zero-shot debiasing capability. Our SLM achieves LLM-level performance on NER tasks, surpassing traditional NER methods, while significantly reducing computational resource consumption compared to LLMs, thereby striking a outstanding balance. VersusDebias is able to debias any models without retraining, avoiding computational overhead. Moreover, as a universal plug-in module, VersusDebias is equipped with ComfyUI API (Comfyanonymous 2024) and Websocket, enabling it to adapt to thousands of models compatible with ComfyUI by simply replacing the workflow. This structure, with the adaptive capabilities provided by generative adversary, enables VersusDebias to debias any T2I models.
@@
To evaluate the effectiveness of our work, we compare VersusDebias with other methods from two perspectives. First, we compare the demographic bias of three general models with and without VersusDebias. Second, we compare VersusDebias with existing debiasing methods. All comparisons consider both zero-shot and few-shot scenarios with multiple evaluation datasets. The results demonstrate that VersusDebias effectively reduces biases in T2I models and significantly outperforms other methods.
@@
  • We propose VersusDebias, a novel and universal framework to efficently decrease demographic biases in T2I models, which ensure fairness of the generated images.
  • VersusDebias integrates numerous advanced technologies, including MLLM, SLM, prompt engineering, and generative adversary.
  • VersusDebias implements zero-shot debiasing via the ComfyUI API and is transferable without retraining, showcasing its universality and promoting practical application of debiasing.
@@ 2 Related Works
Over the last few years, T2I models like Stable Diffusion and DALL-E have seen increasing adoption (Ramesh et al. 2022; Esser et al. 2024). Many T2I models are able to generate high-quality images efficiently. Nonetheless, several studies prove that these models inherit biases from their training datasets (Chinchure et al. 2023; Wan et al. 2024). Recent research classifies these biases into implicit generative bias and explicit generative bias (Luo et al. 2024a), corresponding to the sociological concepts of implicit bias (Pritlove et al. 2019) and explicit bias (Fridell 2013). Implicit generative bias refers to the phenomenon where, without specific instructions on protected attributes, T2I models tend to generate images that do not consist of demographic facts. For instance, when asked to generate ”a photo of a surgeon”, models tend to produce images featuring male surgeons. Explicit generative bias is a specific type of hallucination, referring to the phenomenon where T2I models tend to generate images that do not consist with prompts containing specific protected attributes. For example, when asked to generate ”a photo of a rich black person” models may sometimes fail to correctly generate an image featuring a black person. In our work, we primarily address the issue of implicit generative bias and avoid the influence of explicit generative bias which disturbs existing methods.
@@
Recent works attempts to debias T2I models at different levels, such as diffusion-level (Shrestha et al. 2024) and prompt-level (Clemmer, Ding, and Feng 2024). Diffusion-level methods interfere with the image generation process, which can lead to uncontrollable side effects on the quality and style of the final image. Moreover, these methods can only be applied to diffusion-based models and are not suitable for transformer-based models like DALL-E3 (Betker et al. 2023). In contrast, debiasing based on prompt engineering, as a post-processing method, is more adjustable and more broadly applicable with less impact on image quality. These methods typically performs NER to identify the bias-inducing parts of the prompt and modifies them according to predefined proportions. However, these methods also face serious challenges. First, traditional prompt engineering methods rely on the assumption that the model correctly generate images based on the prompt, which highlights the problem of hallucinations, especially those caused by explicit generative bias. Second, prompt engineering might have unintended side effects on unrelated attributes. For example, using SDXL-Turbo (Sauer et al. 2023), the proportion of women generated with the prompt ”a photo of a South Asian tennis player” is significantly lower than that generated with the prompt ”a photo of a tennis player” (Luo et al. 2024b). Lastly, the severity of these issues varies across different models, meaning that traditional methods, even if they address the first two issues, can only be applied to specific models, greatly limiting their application. In contrast, our method employs an innovative AG module that addresses these problems, providing unprecedented practicality.
Figure 2: Pipeline of VersusDebias. The yellow boxes represent the components of the Discriminator. The three gray rectangles distinguish the three main stages. The blue rectangles represent data that change with the T2I model, while the red rectangles represent data that is independent of the T2I model.
@@
Due to the tremendous success of ChatGPT-4 (Achiam et al. 2023), LLMs, especially MLLMs, garner significant attention and rapidly evolve. The improved accuracy of MLLMs addresses the significant shortcomings of traditional alignment models like CLIP (Radford et al. 2021) and BLIP-2 (Li et al. 2023) in recognizing race and age (Chinchure et al. 2023), providing essential support for the implementation of the alignment module in our discriminator. On the other hand, the development of edge AI attracts more researchers to focus on SLMs (Bai et al. 2023; Zhang et al. 2024), which retain most of the original performance of LLMs while significantly reducing computational resource consumption. Compared to traditional NER methods, LLM-based methods exhibit higher accuracy and zero-shot capabilities (Clemmer, Ding, and Feng 2024), which also indicates the potential of SLMs in NER. In VersusDebias, we utilize a recent SLM and a MLLM to achieve outstanding debiasing performance and accuracy with minimal computational resources.
@@ 3.1 Overview
In this section, we introduce the pipeline of VersusDebias. A brief overview is shown in Figure 2. VersusDebias consists of a AG module, a IG module, and their datasets.
To debias a model, VersusDebias first uses prompts without protected attributes from the dataset to drive the generator to generate images, which are then conveyed into the discriminator. The discriminator employs a MLLM to perform semantic alignment on these images, extracting information on protected attributes. Due to the lack of statistic, we only consider the debiasing of gender, race, and age of the depicted persons. The alignment results are averaged to determine the generative demographic proportions of protected attributes for each prompt. The discriminator compares the proportions of each gender, race, and age group to the ground truth and modifies the prompts by adding the ignored protected attributes, which ensures that the generator produces more images featuring these ignored social groups in the next iteration. The added protected attributes are recorded in the protected attribute array corresponding to each prompt. The newly images generated from modified prompts are then conveyed into the discriminator again, and the process is repeated until the specified number of epochs.
The discriminator then evaluates the generation results for each prompt at each epoch, calculating the cosine similarity between the generative proportions and the ground truth to quantify the conformity between them. The protected attribute array with the highest cosine similarity for each prompt across different epochs is then combined into a single dictionary, referred to as the best results, and conveyed to the IG module.
For generating debiased images with the IG module, the executor uses the fine-tuned SLM to do NER accurately by extracting the potentially bias-inducing part from the input user prompt. It then selects an appropriate protected attribute based on the best results and adds it to the original prompt. Finally, the generator uses the processed prompt to generate debiased images.
@@ 3.2 Generator
As a universal debiasing framework based on prompt engineering, VersusDebias can use any model in its generator, including diffusion-based models (Nichol et al. 2021; Sauer et al. 2023; Esser et al. 2024; Lin, Wang, and Yang 2024; Song, Sun, and Yin 2024) and transformer-based models (Dayma et al. 2021; Ding et al. 2022; Ramesh et al. 2022; Chen et al. 2023). To lower the barrier of VersusDebias and attract more researchers to debiasing studies, we develop an interface with ComfyUI, which is a widely-used GUI for thousands of T2I models based on Stable Diffusion. ComfyUI only requires changes to the input workflow file to adapt to various models and is equipped with a comprehensive API for external program calls. Additionally, due to its openness and flexibility, some models that use new architectures, such as Playground (Li et al. 2024), PixArt (Chen et al. 2024b), and Stable Cascade, are also compatible with ComfyUI by simply installing specific extension nodes. In terms of communication, VersusDebias uses Websocket to monitor the status of ComfyUI and convey instructions. This approach not only achieves a fully automated debiasing process but also allows both systems to operate in independent environments, ensuring optimal compatibility.
In IG module, the generator uses the prompt processed by the executor to generate images based on user input. In AG module, the generator uses prompts from our dataset. The creation process for the dataset is as follows. Based on the 607 detailed occupations classified by the standard occupational classification of the U.S. (BLS 2024a), we consolidate occupations that are typically not distinguished in everyday contexts, resulting in 103 occupations. We also add descriptions of image quality and character states with ChatGPT-4o, generating 20 prompts for each occupation and totaling 2,060 prompts. The detailed process is provided in the supplementary material.
@@ 3.3 Discriminator
As the core part of the AG module, the discriminator consists of three components: the aligner, the array editor, and the evaluator. In the following sections, we provide a detailed explanation of the implementation of them.
@@
The aligner’s function is to align protected attributes with individuals in the images and compute the average of the responses of each occupation. We use the fine-tuned InternVL-4B-1.5 (Chen et al. 2024c) from BIGbench for alignment, which is trained with the dataset from FairFace (Karkkainen and Joo 2021) and reportedly achieves an accuracy of 97.93%. Using race alignment as an instance, the aligner asks the model ”Tell me the race of the main person in the image, White, Black, East Asian, South Asian, or unknown”, and store the response except ”unknown”. The aligner clean the history and tries again if ”unknown”, and skip the image if still ”unknown” persists, assuming the T2I model failed to generate human image. We do a test with results shown in the supplementary material, indicating that this model meeting the requirements of VersusDebias.
@@
The function of the array editor is comparing the generative demographic proportions to the ground truth and adding the underrepresented protected attributes to the prompt. To achieve this, the array editor generates a new dictionary by subtracting the ground truth values from the actual generative proportions. The keys in this dictionary are the occupations, with negative values indicating that the generated data is below the actual data and positive values indicating the opposite. For all protected attributes with negative differences, the array editor adds these underrepresented attributes to the attribute array. The length of the attribute array $L$ can be adjusted based on the requirements for speed and accuracy, but it should generally be maintained at 4 times the number of epochs. This is because there are a total of nine protected attributes, and assuming an equal distribution of positive and negative differences, approximately 4.5 attributes are added each time. Allowing for a buffer, the multiplier is set to 4 as a floor value, and $L$ is at least 5 as a ceiling value. The array editor performs this operation once per epoch until the limit is reached. A detailed algorithm is shown in Algorithm 1.
@@
The ground truth uses the same occupational classification as the prompt set. For race, we employ the four-category classification from FairFace (Karkkainen and Joo 2021), distinguishing race into White, Black, East Asian, and South Asian. For age, we classify it into three stages: young (0-30), middle-aged (31-60), and elderly (60+), based on common age classifications. For gender, due to the cognitive limitations and the diverse of sexual minority identities, we classify gender into male and female, which limits our method to debias for sexual minorities. For race proportions, we use the global racial population data from the UN (UNDESA 2022), as VersusDebias is designed for worldwide users. For gender and age, the proportions for practitioners for a occupation are calculated by dividing the number of individuals of a specific group by the total number of individuals within all detailed occupation belonging to the occupation, based on the data from the BLS (BLS 2024b, c), due to a lack of global data on occupation practitioners and this data also following the SOC.
@@
where $S_{k}$ is the cosine similarity for the protected attribute $k$ of a occupation, $p_{k}i$ and $q_{k}i$ are the generative proportion and real demographic proportion of the $i$ attribute, and $n_{k}$ is the total number of the attributes for this protected attribute.
After calculating the cosine similarity for these three attributes separately, the evaluator computes a weighted average to obtain the cumulative cosine similarity. We set the weights for gender and race at 0.35 and the weight for age at 0.3, which reflects the general consensus that gender and racial discrimination is more common and severe. The weights can be adjusted based on different requirements:
@@
where $S_{c}$ is the cumulative cosine similarity for a occupation, $w_{k}$ is the weight for the protected attribute $k$.
Finally, the evaluator selects the best-performing debiased attribute arrays based on the cumulative cosine similarity and combines them into a single dataset called ”best results”.